# Non-Convex Self-Concordant Functions: Practical Algorithms and Complexity Analysis

### v1

@@ 1 Introduction
Newton’s method is one of the cornerstones of optimization, dating back to the early development of iterative methods. In convex optimization, it is known for its local quadratic convergence when the objective function is smooth and strongly convex [Wright-2006-Numerical]. Despite its efficiency in the local sense, the Newton method often lacks robustness in non-convex settings: without appropriate safeguards, it converges to saddle points or exhibit global convergence slower than first-order methods [Cartis-2010-Complexity]. To address these limitations, extensive research has focused on enhancing its global behavior through various regularization techniques. Among them, cubic regularization methods [Nesterov-2006-Cubic, Nesterov-2008-Accelerating, Cartis-2011-Adaptive, Cartis-2011-Adaptive2] control step size by penalizing large curvature through a cubic term; trust-region methods [Conn-2000-Trust, Curtis-2017-Trust, Curtis-2018-Concise, Curtis-2023-Worst] restrict each update to a region where the local quadratic model is accurate; and regularized Newton methods [Mishchenko-2023-Regularized, Doikov-2024-Gradient, Gratton-2025-Yet] balance curvature exploitation and stability via adaptive damping terms. Collectively, these approaches have established rigorous global convergence guarantees and strong empirical performance across a broad range of non-convex optimization problems.
@@
However, these approaches suffer from some drawbacks [Mishchenko-2023-Regularized, Gratton-2025-Yet]. A primary one is computational: both cubic regularization methods [Nesterov-2006-Cubic] and trust-region methods [Conn-2000-Trust] require solving a nontrivial subproblem at each iteration that involves cubic or constrained quadratic models rather than updating via a simple Hessian inversion. This additional subproblem limits scalability to high-dimensional or large-scale problems. Moreover, their theoretical guarantees depend on strong smoothness conditions, such as the Lipschitz continuity of the Hessian [Nesterov-2006-Cubic, Nesterov-2008-Accelerating, Cartis-2011-Adaptive, Cartis-2011-Adaptive2, Conn-2000-Trust, Curtis-2017-Trust, Curtis-2023-Worst, Mishchenko-2023-Regularized, Doikov-2024-Gradient, Gratton-2025-Yet], and sometimes even of the gradient [Curtis-2017-Trust, Curtis-2023-Worst, Gratton-2025-Yet]. These conditions are often unrealistic in various applications, where Hessian information may be ill-behaved or expensive to approximate accurately.
@@
In the convex setting, these challenges can be effectively mitigated through the notion of self-concordance [Nesterov-1994-Interior, Nesterov-2018-Lectures], which provides a unified framework for studying second-order methods without strong smoothness. Self-concordant functions may possess unbounded third-order derivatives, yet the growth of derivatives is locally controlled by the curvature encoded in the Hessian. This structure ensures that Newton steps remain well-behaved in regions where strong smoothness fails. A canonical example is the logarithmic barrier function, which is self-concordant but lacks a Lipschitz continuous gradient or Hessian. For such functions, second-order methods (e.g., the damped Newton) achieve global convergence and local quadratic convergence, offering both theoretical elegance and computational practicality. Despite its success in convex optimization, extending self-concordance to non-convex settings remains nontrivial. Although its definition appears superficially independent of convexity, it involves the square root of local curvature terms, which can become ill-defined when the Hessian has negative eigenvalues. One might attempt to circumvent this by taking absolute values before square roots, but this modification imposes restrictive curvature bounds that fail to capture simple non-convex structures; for example, the sigmoid function $x\mapsto 1/(1+e^{-x})$ violates this property. This gap motivates developing a generalized notion of self-concordance suitable for non-convex objectives, preserving its curvature control while accommodating regions of negative curvature.
Related works
@@
The classical notion of self-concordance was introduced by [Nesterov-1994-Interior] as a fundamental tool for analyzing interior-point methods in convex optimization. Since then, this notion was extended in various contexts, including composite optimization problems [Tran-2015-Composite], distributed Newton methods [Zhang-2015-Disco], randomized Newton updates [Lu-2017-Randomized] and stochastic quasi-Newton methods [Gao-2019-Quasi]. In parallel, other works studied stronger variants of this notion [Rodomanov-2021-Greedy, Hanzely-2022-Damped], which impose tighter curvature control to improve numerical stability. A recent line of works have began to adopt the generalized notion of quasi-self-concordance, which encompasses important non-quadratic objectives such as logistic regression [Bach-2010-Self, Tran-2015-Composite, Bach-2014-Adaptivity, Zhang-2023-Online, Lee-2024-Improved, Doikov-2025-Minimizing]. Further generalization was proposed by [Tran-2019-Self], who introduced the notion of generalized self-concordance. Building upon this formulation, subsequent works developed and analyzed Frank-Wolfe methods, demonstrating the versatility of self-concordance-based curvature control beyond the standard convex setting [Dvurechensky-2020-Self, Carderera-2021-Simple, Dvurechensky-2023-Generalized].
@@
Our proposed algorithms can also be viewed as second-order methods for non-convex optimization. In this setting, numerous regularization techniques have been developed to ensure global convergence of Newton-type methods. Classical strategies include cubic regularization [Nesterov-2006-Cubic, Nesterov-2008-Accelerating], trust-region frameworks [Conn-2000-Trust], and more recent regularized Newton variants [Mishchenko-2023-Regularized, Doikov-2024-Gradient], all designed to stabilize Newton steps in the presence of non-convex curvature. A major milestone in this line of work is the adaptive cubic regularization method [Cartis-2011-Adaptive, Cartis-2011-Adaptive2], which established a global iteration complexity of $O(\epsilon^{-1.5})$ for minimizing functions with Lipschitz continuous Hessian. This framework has inspired extensive research aimed at improving regularization strategies, reducing computational cost, and refining complexity guarantees [Curtis-2017-Trust, Curtis-2018-Concise, Curtis-2021-Trust, Curtis-2023-Worst, Dussault-2024-Scalable, Gratton-2025-Yet]. To enhance computational efficiency, several works have exploited Hessian-vector products as major oracles, achieving convergence rates on the order of $O(\epsilon^{-7/4})$ while avoiding explicit Hessian computations [Agarwal-2017-Finding, Carmon-2018-Accelerated]. A complementary line of research [Mccormick-1977-Modification, More-1979-Use, Goldfarb-1980-Use] has leveraged negative curvature directions to escape saddle points and approach approximate second-order stationary points. Building on these ideas, [Royer-2018-Complexity, Royer-2020-Newton] proposed hybrid line-search algorithms that alternate between Newton updates and negative-curvature steps identified through the Lanczos method [Kuczynski-1992-Estimating], offering a robust trade-off between curvature exploitation and global progress.
@@
    We introduce two function classes – weakly self-concordant and $F$-based self-concordant functions – that extend the classical notion of self-concordance to non-convex settings. These classes preserve key curvature control properties while encompassing many objectives arising in machine learning.
@@
    We propose regularized Newton and adaptive regularization methods and prove their global and local convergence guarantees without the gradient or Hessian Lipschitz continuity. We extend the adaptive regularization methods [Cartis-2011-Adaptive, Cartis-2011-Adaptive2, Cartis-2022-Evaluation] to constrained problems with self-concordant barriers.
@@
In Section 2, we define non-convex self-concordant functions, and provide three examples of their applicability. In Section 3, we describe our methods and present convergence results. In Section 4, we provide several lemmas and prove our convergence results. In Section 5, we present empirical results complementing our theoretical results. We conclude in Section 6.
2 Non-Convex Self-Concordant Functions
Letting $f:\mathbb{R}^{n}\rightarrow\mathbb{R}\cup\{+\infty\}$ where $\textbf{dom}(f)=\{x\in\mathbb{R}^{n}\mid f(x)<+\infty\}$, the problem of interest is given by
@@ 2 Non-Convex Self-Concordant Functions
In the convex setting, the self-concordant function is defined in [Nesterov-2018-Lectures, Definition 5.1.1]. For the sake of completeness, we summarize it as follows,
@@
A closed and convex function $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ is $\kappa$-self-concordant if $\textbf{dom}(f)$ is open, $f$ is three times continuously differentiable on $\textbf{dom}(f)$, and $\nabla^{3}f(x)[h,h,h]\leq 2\kappa\big(\nabla^{2}f(x)[h,h]\big)^{3/2}$ for all $x\in\textbf{dom}(f)$ and $h\in\mathbb{R}^{n}$.
We next extend this notion to non-convex settings.
@@
A function $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ is $(\kappa,\ell)$-weakly self-concordant if the function $f(x)+\frac{\ell}{2}\|x\|^{2}$ is $\kappa$-self-concordant.
The weakly self-concordant functions form a subclass of weakly convex functions [Davis-2019-Stochastic]. Indeed, $f$ is $(\kappa,\ell)$-weakly self-concordant if and only if $f(x)+\frac{\ell}{2}\|x\|^{2}$ is closed and convex, and for all $x\in\textbf{dom}(f)$ and $h\in\mathbb{R}^{n}$, we have
$\nabla^{3}f(x)[h,h,h]=\nabla^{3}\left(f+\tfrac{\ell}{2}\|\cdot\|^{2}\right)(x)[h,h,h]\leq 2\kappa\left(\nabla^{2}f(x)[h,h]+\ell\|h\|^{2}\right)^{3/2}.$
In particular, if $f$ is $\ell$-weakly convex and satisfies $|\nabla^{3}f(x)[h,h,h]|\leq m\|h\|^{3}$, then $f$ is $(\frac{m}{2a^{3/2}},\ell+a)$-weakly self-concordant for any $a>0$.
@@
Given a reference function $F:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$, we say that $f$ is $F$-based $\kappa$-self-concordant if $\textbf{dom}(f)\subseteq\textbf{dom}(F)$ and $f+F$ is $\kappa$-self-concordant.
Classical and weakly self-concordant functions are two special instances of general $F$-based self-concordant functions. This formulation unifies analysis across different settings and enjoys convenient closure properties.
@@
By definition, we have $f_{i}+F_{i}$ is $\kappa_{i}$-self-concordant for $i=1,2$. Then, [Nesterov-2018-Lectures, Theorem 5.1.1] shows that $\alpha_{1}(f_{1}+F_{1})+\alpha_{2}(f_{2}+F_{2})$ is $\max(\frac{\kappa_{1}}{\sqrt{\alpha_{1}}},\frac{\kappa_{2}}{\sqrt{\alpha_{2}}})$-self-concordant. This implies the desired result. ∎
@@ 2.1 Examples
In what follows, we present three representative examples of $F$-based self-concordant functions with full details.
Generalized phase retrieval
As shown in [Sun-2018-Geometric], the loss function $f:\mathbb{R}^{2n}\to\mathbb{R}$ is defined as follows ($a_{k}^{*}$ represents the conjugate transpose of $a_{k}$),
$f(z^{R},z^{I})=\tfrac{1}{2m}\sum_{k=1}^{m}(|a_{k}^{*}z|^{2}-y_{k}^{2})^{2}\textnormal{ for }a_{k}=\tfrac{1}{\sqrt{2}}(a_{k}^{R}+a_{k}^{I}\sqrt{-1})\in\mathbb{C}^{n},\ z=z^{R}+z^{I}\sqrt{-1}.$
This is a $(\kappa,\ell)$-weakly self-concordant function for some $\kappa>0$ and $\ell>0$. Indeed, we can rewrite the loss function as
$f(z^{R},z^{I})=\tfrac{1}{2m}\sum_{k=1}^{m}\left(\tfrac{1}{2}((a_{k}^{R})^{\top}z^{R}+(a_{k}^{I})^{\top}z^{I})^{2}+\tfrac{1}{2}((a_{k}^{R})^{\top}z^{I}-(a_{k}^{I})^{\top}z^{R})^{2}-y_{k}^{2}\right)^{2}.$
By Proposition 2.4, we claim that $g(x)=((c_{1}^{\top}x)^{2}+(c_{2}^{\top}x)^{2})^{2}$ is weakly self-concordant. Indeed, we denote $c_{i}^{\top}x$ by $x_{i}$ and $c_{i}^{\top}h$ by $h_{i}$ for $i=1,2$. Then, we have $\nabla^{2}g(x)[h,h]=8(x_{1}h_{1}+x_{2}h_{2})^{2}+4(x_{1}^{2}+x_{2}^{2})(h_{1}^{2}+h_{2}^{2})$. For some $\ell\geq\|c_{1}\|^{2}+\|c_{2}\|^{2}$, we have
$(\nabla^{2}g(x)[h,h]+\ell\|h\|^{2})^{\frac{3}{2}}\geq\sum_{i,j\in\{1,2\}}(|x_{i}|^{3}|h_{j}|^{3}+|h_{i}|^{3}+|h_{j}|^{3}).$
We also have
$\nabla^{3}g(x)[h,h,h]=24(x_{1}h_{1}+x_{2}h_{2})(h_{1}^{2}+h_{2}^{2})\leq 8\sum_{i,j\in\{1,2\}}(|x_{i}|^{3}|h_{j}|^{3}+|h_{i}|^{3}+|h_{j}|^{3})$
Putting these pieces together yields that $g$ is $(4,\ell)$-weakly self-concordant.
The above results can be generalized; indeed, any polynomial functions are $F$-based self-concordant for some functions $F$.
@@
Let $p\geq 2$ be integer-valued. Suppose that $f:\mathbb{R}^{n}\to\mathbb{R}$ is a multivariate polynomial with degree $\deg(f)\leq 2p$. Then, $F(x)=(\|x\|^{2}+1)^{p}$ is convex and there exists $m\geq 0$ such that $f$ is $mF$-based 11-self-concordant.
@@
Since $\deg(f)\leq 2p$, each entry of $\nabla^{2}f(x)$ (resp., $\nabla^{3}f(x)$) is a polynomial of degree at most 2p22p-2 (resp., at most 2p32p-3). There exist $c_{1},c_{2}\geq 0$ such that, for all $x,h\in\mathbb{R}^{n}$, we have $|\nabla^{2}f(x)[h,h]|\leq c_{1}(1+\|x\|^{2})^{p-1}\|h\|^{2}$ and $|\nabla^{3}f(x)[h,h,h]|\leq c_{2}(1+\|x\|^{2})^{p-3/2}\|h\|^{3}$. In addition, we have $|\nabla^{3}F(x)[h,h,h]|\leq c_{3}(1+\|x\|^{2})^{p-\frac{3}{2}}\|h\|^{3}$.
We define $\varphi(r)=(r+1)^{p}$. Then, we have $\nabla F(x)=2\varphi^{\prime}(\|x\|^{2})x$, $\nabla^{2}F(x)=2\varphi^{\prime}(\|x\|^{2})I+4\varphi^{\prime\prime}(\|x\|^{2})xx^{\top}$. For all $h\in\mathbb{R}^{n}$, we have
@@
Thus, $F$ is convex. Define $g=f+mF$, we have $\nabla^{2}g(x)[h,h]\geq(2pm-c_{1})(\|x\|^{2}+1)^{p-1}\|h\|^{2}$. Moreover, $|\nabla^{3}g(x)[h,h,h]|\leq(c_{2}+mc_{3})(\|x\|^{2}+1)^{p-3/2}\|h\|^{3}$. Thus, there exists $m\geq 0$ such that $g=f+mF$ is $1$-self-concordant. ∎
@@
We consider the constrained polynomial optimization problem $\min_{x\in C}f(x)$, where $f$ is a polynomial, and suppose there exists a self-concordant function $g:\mathbb{R}^{n}\to\mathbb{R}$ such that $\textbf{dom}(g)=C$. By Proposition 2.5, we have $f+F$ is self-concordant on $\mathbb{R}^{n}$. Thus, we have $f+F+g$ is self-concordant on $C$ and $f\mid_{C}$ is $(F+g)$-based self-concordant. This perspective provides numerous examples of $F$-based self-concordant functions, such as linear regression, matrix factorization, sensor network localization [Tang-2024-Riemannian], and $D$-optimal design [Todd-2016-Minimum], without applying the logarithm in the objective and with the sum of decision variables being less than $1$ rather than equal to $1$.

### v2

@@ 1 Introduction
Newton’s method is one of the cornerstones of optimization, dating back to the early development of iterative methods. In convex optimization, it is known for its local quadratic convergence when the objective function is smooth and strongly convex [60]. However, Newton’s method is not suitable in non-convex settings without appropriate safeguards, as it may converge to saddle points or exhibit slower global convergence than first-order methods [10]. To address these limitations, extensive research has focused on enhancing its global behavior through various regularization techniques. Among them, cubic regularization methods [31, 56, 57, 11, 12] control the search direction size by penalizing large curvature through a cubic term; trust-region methods [15, 18, 16, 19] restrict each update to a region where the local quadratic model is accurate; and regularized (also referred to as “damped”) Newton’s methods [40, 46, 74, 63, 52, 22, 30] balance curvature exploitation and stability via adaptive damping terms. Collectively, these approaches have established rigorous global convergence guarantees and strong empirical performance across a broad range of non-convex optimization problems.
@@
However, these approaches suffer from some drawbacks [52, 30]. A primary one is implementation: both cubic regularization methods [56] and trust-region methods [15] require solving a nontrivial subproblem at each iteration that involves cubic or constrained quadratic models rather than a simple Hessian inversion. While efficient strategies exist for these subproblems (see, e.g., [4, 23, 29, 3]), they nevertheless introduce additional implementation complexity compared to a simple regularized Newton update, and their efficient deployment in high-dimensional settings remains an active area of research. Moreover, the theoretical guarantees of cubic regularization methods and trust region methods depend on strong smoothness conditions, such as the Lipschitz continuity of the Hessian [56, 57, 63, 11, 12, 15, 18, 19, 52, 22, 30], and sometimes even of the gradient [18, 19, 30]. These conditions are often unrealistic in various applications.
@@
In the convex setting, the notion of self-concordance [55, 58], (see Definition 2.1 below), provides a unified framework for studying second-order methods without Hessian Lipschitz continuity. More specifically, a convex and three times continuously differentiable function $f:\textbf{dom}(f)\to\mathbb{R}$ is defined to be $\kappa$-self-concordant if for all $x\in\textbf{dom}(f)$, $\nabla^{3}f(x)[h,h,h]\leq 2\kappa(\nabla^{2}f(x)[h,h])^{1.5}$. In particular, self-concordant functions may possess unbounded third-order derivatives, yet the growth of derivatives is locally controlled by the curvature encoded in the Hessian. This structure ensures that Newton steps remain well-behaved in regions where Hessian Lipschitz continuity fails. A canonical example is the logarithmic barrier function, which is self-concordant but lacks a Lipschitz continuous gradient or Hessian. For such functions, second-order methods (e.g., damped Newton) achieve global convergence and local quadratic convergence, offering both theoretical guarantees and computational practicality. Unfortunately, extending self-concordance to non-convex settings remains nontrivial since its definition involves the square root of a local curvature term, which can become ill-defined when the Hessian has negative eigenvalues. One might attempt to define self-concordance as $\nabla^{3}f(x)[h,h,h]\leq 2\kappa(|\nabla^{2}f(x)[h,h]|)^{1.5}$ when $f$ is potentially non-convex, but this modification imposes restrictive curvature bounds that fail to capture simple non-convex structures; for example, the sigmoid function $x\mapsto 1/(1+e^{-x})$ violates this property. This gap motivated our development of (i) a generalized notion of self-concordance suitable for non-convex objective functions that preserves its curvature control while accommodating regions of negative curvature, and (ii) the development and analysis of second-order algorithms that exploit this property.
Related work on self-concordance
@@
The classical notion of self-concordance was introduced by [55] as a fundamental tool for analyzing interior-point methods in convex optimization [79, 59, 32]. Since then, this notion has been extended to composite optimization problems [75], distributed Newton methods [84], randomized Newton updates [44] and stochastic quasi-Newton methods [27]. In parallel, other work has studied stronger variants of this notion [65, 33], which impose tighter curvature control to improve numerical stability. Another line of work has adopted the notion of quasi-self-concordance, which encompasses important objectives such as logistic regression [5, 75, 6]. Finally, in [72], the notion of generalized self-concordance was introduced. Building upon this, Frank-Wolfe methods have been developed that demonstrate the versatility of self-concordance-based curvature control beyond the standard convex setting [24, 8, 25]. We note that all of the above extensions of self-concordance apply only to convex functions.
Related work on regularized methods for non-convex optimization
@@
Numerous regularization techniques have been developed to ensure global convergence of Newton-type methods that are applied to minimizing non-convex functions. Such strategies include cubic regularization [56, 57], trust-region  [15], and recent regularized Newton variants [52, 22, 85, 30], all designed to stabilize Newton steps in the presence of non-convex curvature. A major milestone in this line of work is the adaptive cubic regularization method [11, 12], which established a global iteration complexity of $O(\epsilon^{-1.5})$ for minimizing functions with Lipschitz continuous Hessians. This framework has inspired extensive research aimed at improving regularization strategies, reducing computational cost, and refining complexity guarantees [18, 16, 17, 19, 23, 30, 81]. Adaptive cubic regularization has also been extended to constrained non-convex settings, including problems with convex constraints [14] and those involving conic constraints equipped with logarithmically homogeneous self-concordant barriers [26]. To enhance computational efficiency, several works have exploited Hessian-vector products as major oracles, achieving convergence rates on the order of $O(\epsilon^{-7/4})$ while avoiding explicit Hessian computations [4, 9]. A complementary line of research [51, 53, 28] has leveraged negative curvature directions to escape saddle points and approach approximate second-order stationary points. Building on these ideas, [67, 66, 34] proposed hybrid line-search algorithms that alternate between Newton updates and negative-curvature steps identified through the Lanczos method [38], offering a robust trade-off between curvature exploitation and global progress.
@@
    We introduce two function classes – weakly self-concordant and $F$-based self-concordant functions – that extend the classical notion of self-concordance to non-convex settings. These classes preserve key curvature properties while encompassing many objectives arising in machine learning.
@@
    We propose regularized Newton and adaptive regularization methods that make use of our weakly and $F$-based self-concordant notions and prove their global and local convergence without requiring gradient or Hessian Lipschitz continuity, and any line-search. We extend the adaptive regularization methods [11, 12, 13] to constrained problems with self-concordant barriers, where the regularization parameter is tuned automatically.
@@
In Section 2, we define non-convex self-concordant functions, and provide three examples of their applicability. In Section 3, we describe our methods and present convergence results. In Section 4, we present empirical results complementing our theoretical results, and offer some conclusions in Section 5. In an Appendix we present some additional results on weakly self-concordant functions, including an inexact proximal point method for minimizing them.
2 Non-Convex Weakly and $F$-based Self-Concordant Functions
Letting $f:\mathbb{R}^{n}\rightarrow\mathbb{R}\cup\{+\infty\}$ be three-times continuously differentiable on $\textbf{dom}(f)=\{x\in\mathbb{R}^{n}\mid f(x)<+\infty\}$, we consider the problem
@@ 2 Non-Convex Weakly and $F$-based Self-Concordant Functions
[58, Definition 5.1.1] Let $\kappa>0$. A closed and convex function $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ is $\kappa$-self-concordant if $\textbf{dom}(f)$ is open, $f$ is three times continuously differentiable on $\textbf{dom}(f)$, and $\nabla^{3}f(x)[h,h,h]\leq 2\kappa\big(\nabla^{2}f(x)[h,h]\big)^{3/2}$ for all $x\in\textbf{dom}(f)$ and $h\in\mathbb{R}^{n}$.
We now extend this notion to non-convex settings.
@@
A function $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ is $(\kappa,\ell)$-weakly self-concordant if the function $g(x)=f(x)+\frac{\ell}{2}\|x\|^{2}$ is $\kappa$-self-concordant and satisfies $\nabla^{2}g(x)\succ 0$ for all $x\in\textbf{dom}(f)$.
Weakly self-concordant functions form a subclass of weakly convex functions [21]. Indeed, $f$ is $(\kappa,\ell)$-weakly self-concordant if and only if $f(x)+\frac{\ell}{2}\|x\|^{2}$ is closed, convex, with a Hessian that is non-degenerate, and
$\nabla^{3}f(x)[h,h,h]=\nabla^{3}\left(f+\tfrac{\ell}{2}\|\cdot\|^{2}\right)(x)[h,h,h]\leq 2\kappa\left(\nabla^{2}f(x)[h,h]+\ell\|h\|^{2}\right)^{3/2},$
for all $x\in\textbf{dom}(f)$ and $h\in\mathbb{R}^{n}$. In particular, if $f$ is $\ell$-weakly convex and satisfies $|\nabla^{3}f(x)[h,h,h]|\leq m\|h\|^{3}$ for all $x\in\textbf{dom}(f)$, then $f$ is $(\frac{m}{2a^{3/2}},\ell+a)$-weakly self-concordant for any $a>0$.
The requirement of $\nabla^{2}g(x)\succ 0$ is technical and can be circumvented by choosing a larger $\ell$, assuming that $f$ is weakly convex. In fact, when implementing variants of Newton’s method for convex self-concordant functions, the objective is also assumed to be strictly convex [58].
@@
Given a base function $F:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$, we say that $f$ is $F$-based $\kappa$-self-concordant if $\textbf{dom}(f)\subseteq\textbf{dom}(F)$ and $g=f+F$ is $\kappa$-self-concordant and satisfies $\nabla^{2}g(x)\succ 0$ for all $x\in\textbf{dom}(f)$.
Again, the requirement of $\nabla^{2}g(x)\succ 0$ is technical and can be circumvented by adding a quadratic term to $F$. Classical strictly convex self-concordant functions are 0-based self-concordant. Moreover, $(\kappa,\ell)$-weakly self-concordance is equivalent to $\frac{\ell}{2}\|\cdot\|^{2}$-based $\kappa$-self-concordance. This formulation of $F$-based self-concordance unifies analysis across different settings and enjoys convenient closure properties.
@@
Proof Since $f_{i}+F_{i}$ is $\kappa_{i}$-self-concordant for $i=1,2$,  [58, Theorem 5.1.1] shows that $\alpha_{1}(f_{1}+F_{1})+\alpha_{2}(f_{2}+F_{2})$ is $\max(\frac{\kappa_{1}}{\sqrt{\alpha_{1}}},\frac{\kappa_{2}}{\sqrt{\alpha_{2}}})$-self-concordant. This implies the desired result. $\square$
@@ 2.1 Examples
We now present three representative examples of $F$-based self-concordant functions with full details.
Generalized Phase Retrieval (GPR)
As in [71], the GPR loss function $f:\mathbb{R}^{2n}\to\mathbb{R}$ is defined as
$f(z^{R},z^{I})=\tfrac{1}{2m}\sum_{k=1}^{m}(|a_{k}^{*}z|^{2}-y_{k}^{2})^{2},\quad\textnormal{ where }a_{k}=\tfrac{1}{\sqrt{2}}(a_{k}^{R}+a_{k}^{I}\sqrt{-1})\in\mathbb{C}^{n},\ z=z^{R}+z^{I}\sqrt{-1},$
and $a_{k}^{*}$ represents the conjugate transpose of $a_{k}$. $f(z^{R},z^{I})$ is a $(\kappa,\ell)$-weakly self-concordant function for some $\kappa>0$ and $\ell>0$. Indeed, we can write $f(z^{R},z^{I})$ as
$f(z^{R},z^{I})=f(x)=\tfrac{1}{2m}\sum_{k=1}^{m}\left(\tfrac{1}{2}((c(1)_{k}^{\top}x)^{2}+(c(2)_{k}^{\top}x)^{2})-y_{k}^{2}\right)^{2},$
where $x^{\top}=((z^{R})^{\top},(z^{I})^{\top})$, $c(1)_{k}^{\top}=((a_{k}^{R})^{\top},(a_{k}^{I})^{\top})$ and $c(2)_{k}^{\top}=(-(a_{k}^{I})^{\top},(a_{k}^{R})^{\top})$.
By Proposition 2.4, to prove that $f(z^{R},z^{I})=f(x)$ is $F$-based or weakly self-concordant we only need to show this for each term in the above summation, and in particular for $g(x)=((\bar{x}_{1})^{2}+(\bar{x}_{2})^{2})^{2}$, where for the $k$-th term $\bar{x_{i}}\equiv(c(i)_{k}^{\top}x)$ for $i=1,2$. Defining $\bar{h_{i}}\equiv(c(i)_{k}^{\top}h)$ for $i=1,2$, we have that
$\nabla^{2}g(x)[h,h]=8(\bar{x}_{1}\bar{h}_{1}+\bar{x}_{2}\bar{h}_{2})^{2}+4(\bar{x}_{1}^{2}+\bar{x}_{2}^{2})(\bar{h}_{1}^{2}+\bar{h}_{2}^{2}),$
and for $\ell\geq\|c(1)_{k}\|^{2}+\|c(2)_{k}\|^{2}$, that
$(\nabla^{2}g(x)[h,h]+\ell\|h\|^{2})^{\frac{3}{2}}\geq(\bar{h}_{1}^{2}+\bar{h}_{2}^{2})^{3/2}\left(4(\bar{x}_{1}^{2}+\bar{x}_{2}^{2})+1\right)^{3/2},$
$\nabla^{3}g(x)[h,h,h]=24(\bar{x}_{1}\bar{h}_{1}+\bar{x}_{2}\bar{h}_{2})(\bar{h}_{1}^{2}+\bar{h}_{2}^{2})\leq 24\sqrt{\bar{x}_{1}^{2}+\bar{x}_{2}^{2}}\;(\bar{h}_{1}^{2}+\bar{h}_{2}^{2})^{3/2}.$
Hence, $g$ is $(3,\ell)$-weakly self-concordant.
Polynomial Optimization Problems (POPs)
The above results for GPR can be generalized; indeed, any polynomial function is $F$-based self-concordant, and hence, also $F$-based convex”, as shown below in Proposition 2.5, for some function $F$. However, many polynomial functions (e.g., $x^{4}+3x^{2}$) are not weakly convex and hence not weakly self-concordant. Therefore, it is necessary to incorporate the concept of $F$-based self-concordance into our analysis.
@@
Let $p\geq 2$ be an integer. Suppose that $f:\mathbb{R}^{n}\to\mathbb{R}$ is a multivariate polynomial with degree $\deg(f)\leq 2p$. Then, $F(x)=(\|x\|^{2}+1)^{p}$ is convex and there exists $m\geq 0$ such that $f$ is $mF$-based 11-self-concordant.
@@
Proof Since $\deg(f)\leq 2p$ and $\deg(F)=2p$, each term of $\nabla^{2}f(x)$, $\nabla^{3}f(x)$, and $\nabla^{3}F(x)$, is respectively, a polynomial of degree at most 2p22p-2, 2p32p-3, and exactly 2p32p-3. There exist $c_{1},c_{2}\geq 0$ and $c_{3}>0$ such that, for all $x,h\in\mathbb{R}^{n}$, we have $|\nabla^{2}f(x)[h,h]|\leq c_{1}(1+\|x\|^{2})^{p-1}\|h\|^{2}$, $|\nabla^{3}f(x)[h,h,h]|\leq c_{2}(1+\|x\|^{2})^{p-3/2}\|h\|^{3}$, and $|\nabla^{3}F(x)[h,h,h]|\leq c_{3}(1+\|x\|^{2})^{p-\frac{3}{2}}\|h\|^{3}$.
Defining $\varphi(r)=(r+1)^{p}$, we have $\nabla F(x)=2\varphi^{\prime}(\|x\|^{2})x$, $\nabla^{2}F(x)=2\varphi^{\prime}(\|x\|^{2})I+4\varphi^{\prime\prime}(\|x\|^{2})xx^{\top}$, and for all $h\in\mathbb{R}^{n}$,
@@
Thus, $F$ is convex, $\nabla^{2}\left(f+mF\right)(x)[h,h]\geq(2pm-c_{1})(\|x\|^{2}+1)^{p-1}\|h\|^{2}$ and $|\nabla^{3}\left(f+mF\right)(x)[h,h,h]|\leq(c_{2}+mc_{3})(\|x\|^{2}+1)^{p-3/2}\|h\|^{3}$, which show that there exists $m\geq 0$ such that $f+mF$ is $1$-self-concordant. $\square$
@@
We note that there are many constrained POPs, $\min_{x\in C}f(x)$, where $f$ is a polynomial, and there exists a self-concordant barrier function $g:\mathbb{R}^{n}\to\mathbb{R}$ such that $\textbf{dom}(g)=\textbf{int}(C)$. By Proposition 2.5, we have $f+F$ is self-concordant on $\mathbb{R}^{n}$ for some $F:\mathbb{R}^{n}\to\mathbb{R}$. Thus, we have $f+F+g$ is self-concordant on $\textbf{int}(C)$ and $f\mid_{\textbf{int}(C)}$ is $(F+g)$-based self-concordant. This perspective provides numerous examples of $F$-based self-concordant functions, such as matrix factorization $\min_{U,V}\|M-UV^{\top}\|_{F}^{2}$ and sensor network localization [73] $\min_{x_{1},\ldots,x_{n}\in\mathbb{R}^{d}}\sum_{(i,j)\in\mathcal{E}}\bigl(\|x_{i}-x_{j}\|^{2}-d_{ij}^{2}\bigr)^{2}$, where $\mathcal{E}\subseteq\{(i,j):i<j,i,j\in\{1,...,n\}\}$ and $d_{ij}\geq 0$ are constants.